Solve the initial value problem (IVP):


(D^2+4D+4)=9e^x


y(0)=2


y'(0)=1

1
Expert's answer
2022-01-06T17:48:51-0500

"D^2+4D+4=9e^x"

Where "D" is the linear differential operator "\frac{d}{dx}"

Thus we can write the equation as:

"\frac{d^2y}{dx^2}+4\frac{dy}{dx}+4y=9e^x"

The associated auxilliary equation is;

"m^2+4m+4=0\\
(m+2)(m+2)\\
m=-2 \text{ twice}"

Thus the solution of the homogeneous part is

"y_c=(A+Bx)e^{-2x}"

For the particular solution

Consider the solution "y=Ae^x"

"\implies y'=Ae^x,y''=Ae^x"

Substitute this into the equation.

"Ae^x+4Ae^x+4Ae^x=9e^x\\
9Ae^x=9e^x\\
A=1"

Thus the particular solution is;

"y_p=e^x"

Hence, the general solution is

"y=(A+Bx)e^{-2x}+e^x"

Now, consider the initial value given. "y(0)=2,y'(0)=1"

"y(0)=A+1=2\\A=1"

"y'(x)=Be^{-2x}-2(A+Bx)e^{-2x}+e^x\\
y'(0)=B-2A+1=1\\
B=2A=2"

Hence the solution to the IVP is;

"y=(1+2x)e^{-2x}+e^x"